1. A cue ball on a pool table is hit with the cue stick in one corner of the table and rolls to the opposite corner.  The mass of the ball is 0.17 kg and it starts rolling at a speed of 3.6 m/s.  By the time it reaches the opposite corner of the table, the ball has a kinetic energy of 0.65 J.  Determine the following: (a) the initial kinetic energy, (b) the speed of the ball at the opposite corner of the table, and (c) the work done by friction as the ball rolls from one corner to the other.
1
Expert's answer
2022-01-18T13:42:22-0500

"\text{(a)}.\space KE=\dfrac 12mv^2=1.1\text{ J}.\\\space\\
\text{(b)}.\space u=\sqrt{2E/m}=2.8\text{ m/s}.\\\space\\
\text{(c)}.\space W_f=KE-E=1.1-0.65=0.45\text{ J}.\\"